# TyphoFormer: Language-Augmented Transformer for Accurate Typhoon Track Forecasting

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Typhoons and hurricanes pose significant threats to coastal regions, causing devastating damage to infrastructure, ecosystems, and human lives (Ramachandran et al., 2025; Wachnicka et al., 2020; Hao et al., 2024). The southeastern United States, especially Florida, is facing increasingly frequent and intense storms (Oliver-Smith, 2019; Malmstadt et al., 2009). For instance, Hurricane IAN in 2022 resulted in over 150 fatalities and an estimated \$113 billion in economic losses, making it one of the costliest hurricanes in U.S. history (Wang et al., 2024c; Bushong and Welch, 2023). With the rise of sea surface temperatures and the evolution of climate patterns, accurate typhoon trajectory forecasting has become more critical than ever for emergency planning, resource allocation, and early evacuation (Wang et al., 2024a). Reliable predictions of storm movement not only support disaster mitigation efforts but also enable authorities to issue timely warnings, minimizing potential casualties and disruption (Yu et al., 2022). In light of the substantial socio-economic impact of such extreme weather events, there is a growing demand for advanced, data-driven forecasting methods that can complement or enhance traditional meteorological models.
@@
Existing approaches for typhoon track forecasting generally fall into two main categories: (I) time series modeling methods and (II) physics-based numerical models. To address the need for data-driven forecasting, a growing body of work has explored modeling typhoon tracks as temporal sequences using deep learning techniques (Wang et al., 2024b). These methods treat the storm’s historical positions and attributes as a time series and aim to predict its future trajectory based on learned temporal patterns using deep learning-based methods. Early approaches employed recurrent neural network (RNN)-based architectures such as LSTM, GRU, and their variants to capture sequential dependencies in storm movement (Gao et al., 2018; Song et al., 2022; Wei and Cheng, 2020). More recent advances have leveraged Transformer-based architectures, which offer improved long-range modeling capabilities and greater scalability for sequence forecasting (Jiang et al., 2023; Gan et al., 2024). Methods such as InFormer (Zhou et al., 2021) and AutoFormer (Wu et al., 2021) have demonstrated promising results in related time series domains, particularly for long-horizon prediction tasks. While these models excel in learning from past trajectory data, they lack an awareness of external contextual information, such as meteorological semantics or high-level event dynamics, that may further enhance predictive accuracy.
@@
As a basic tool in traditional meteorological analysis, physics-based numerical models are also widely used in operational typhoon forecasting (Iwasaki et al., 1987; Knaff et al., 2007). Numerical models simulate atmospheric dynamics by solving complex partial differential equations derived from physical laws (Bauer et al., 2015), such as the Navier–Stokes Equations (Rozanova et al., 2010) and Boundary Layer Structure Modeling (Kepert, 2010). Prominent forecasting systems, including the the Global Forecast System (GFS) (Halperin et al., 2020) and Climatology and Persistence Model (Knaff et al., 2007), use large-scale initial conditions, such as pressure fields, wind vectors, and sea surface temperatures, to estimate future storm trajectories. While these models offer physical interpretability and have been refined through decades of meteorological research, they require gigantic computational resources (e.g., supercomputing center) and are sensitive to initial condition uncertainties. Furthermore, their performance may degrade in data-sparse regions or during rapidly evolving storm phases, where high-resolution input data are limited or delayed. As a result, complementary deep learning and data-driven techniques have been increasingly explored to address these limitations inherent in traditional approaches.
@@ 2.2. Language Contextual Prompt Generation
To enhance the model’s contextual understanding of meteorological dynamics, we generate natural language descriptions (referred to as semantic prompts) for each time step in the trajectory data. Unlike conventional approaches that rely solely on structured numerical inputs, our method introduces a textual layer of information that captures the same data in a human-readable, semantically coherent format. These prompts are automatically generated using a GPT-4 through API-based batch inference. Given a single line record from the trajectory dataset, such as:
20241010, 0030, L, HU, 27.4N, 82.6W, 100, 958, 180,
170, 110, 220, 60, 60, 70, 90, 30, 30, 30, 30, 20
@@
These natural language prompts encode meteorological knowledge in a compact form that would otherwise require explicit domain understanding. By providing a semantic summary of each state, they serve as high-level guidance to assist the forecasting model in interpreting patterns and trends within the trajectory. All prompts are generated offline using the GPT-4 API and cached for training and evaluation. This step introduces minimal computational overhead and significantly enriches the input representation.
@@ 2.3. Text Embedding and Feature Fusion
Instead of directly concatenating these two modalities, we introduce a Prompt-aware Gating Fusion (PGF) mechanism to adaptively fuse the prompt information into the temporal sequence. First, we compute a condensed prompt representation by averaging the token embeddings:$\mathbf{p}_{\text{mean}}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{p}_{m}$. Then, for each time step $t$, we compute a gated fusion vector:
@@ 3. Experiments
Dataset and Evaluation Metrics. Experiments are carried on the well-known HURDAT2 dataset, which is the official North Atlantic hurricane database maintained by the U.S. National Hurricane Center (NHC). It provides detailed 6-hour records of tropical cyclones, including 22 critical meteorological features. Each typhoon event is recorded as a chronological sequence. We use the records from year 2004 to 2021 for training, while the data samples from 2022 to 2024 are used for testing.
For evaluation metrics, we use the commonly adopted prediction criteria Mean Absolute Error (MAE) and a task-specific criteria–Spherical Distance Error ($\Delta\text{R}$). $\Delta\text{R}$ measures the geospatial distance between the prediction and the ground-truth trajectories, which is calculated as: $\Delta\text{R}=\Delta\text{Re}\cdot\text{arccos}(\sin\varphi_{p}\sin\varphi_{r}+\cos\varphi_{p}\cos\varphi_{r}\cos(\lambda_{p}-\lambda_{r}))$, where $\text{Re}=6371km$ is the Earth’s radius, $(\varphi_{p},\lambda_{p})$ and $(\varphi_{r},\lambda_{r})$ are the predicted / ground-truth latitude and longitude, respectively.
Baseline methods. We compare our TyphoFormer [^2] https://github.com/LabRAI/TyphoFormer against two categories of baselines. The first category is the classical numerical equation-based models, we choose the representative (I) CLIPER (Climatology and Persistence Model) (Knaff et al., 2007), which extrapolates storm trajectories using historical patterns and persistence assumptions. The second category is time-series forecasting models, including recurrent models like (III) LSTM and (IV) GRU, as well as Transformer-based or MLP-based advanced architectures like (V) Informer (Zhou et al., 2021), (VI) Autoformer (Wu et al., 2021), and (VII) TSMixer (Ekambaram et al., 2023). These baselines provide a diverse comparison across classical statistical models and advanced deep learning frameworks.
@@ 3.1. Main Results
The best values are in bold font, the second-best values are indicated with underlining. Models MAE (All) $\Delta$R (km) (All) 6h 12h 18h 24h 6h 12h 18h 24h CLIPER 0.235 0.275 0.310 0.368 34.265 42.205 51.632 58.268 GRU 0.367 0.405 0.493 0.640 50.480 69.397 90.875 103.894 LSTM 0.392 0.431 0.583 0.828 46.096 71.365 95.412 112.663 Informer 0.289 0.318 0.392 0.483 37.592 46.435 56.433 76.684 Autoformer 0.263 0.286 0.357 0.433 39.836 47.183 63.775 70.862 TSMixer 0.214 0.268 0.297 0.353 35.720 45.265 50.330 62.910 TyphoFormer 0.188 0.242 0.261 0.312 31.539 38.084 42.435 49.562 Models MAE (2024) $\Delta$R (km) (2024) 6h 12h 18h 24h 6h 12h 18h 24h CLIPER 0.221 0.269 0.294 0.361 33.915 44.542 48.580 57.642 GRU 0.359 0.394 0.491 0.632 49.674 70.357 92.542 105.114 LSTM 0.388 0.422 0.587 0.835 44.203 69.051 94.555 109.108 Informer 0.290 0.316 0.395 0.487 37.843 48.791 57.445 78.497 Autoformer 0.257 0.284 0.358 0.435 39.128 46.085 65.039 72.827 TSMixer 0.228 0.265 0.305 0.347 35.148 43.935 52.501 63.118 TyphoFormer 0.185 0.241 0.263 0.317 31.274 37.934 42.973 50.881
@@
We report the performance of TyphoFormer and various baseline methods on typhoon track prediction in Table 1, using both Mean Absolute Error (MAE) and great-circle distance error ($\Delta$R) across four prediction horizons (6h, 12h, 18h, 24h). Results are reported on the full dataset (top) and on the 2024 subset (bottom). Overall, TyphoFormer consistently outperforms all baselines across all metrics and time horizons. Compared with the classical statistical model CLIPER and numerical forecast proxy NCEP, our model yields significantly lower MAE and $\Delta$R, especially at longer horizons (e.g., achieving a 24h MAE of 0.312 and $\Delta$R of 49.56 km vs. CLIPER’s 0.368 and 58.27 km, respectively). This demonstrates the advantage of incorporating contextual knowledge through natural language-guided fusion. Among the deep learning baselines, TyphoFormer surpasses both recurrent models (GRU, LSTM) and Transformer-based architectures (Informer, Autoformer, TSMixer). Notably, our model shows the most stable performance across different forecast steps, with minimal degradation over longer horizons. For instance, the 24h $\Delta$R of TyphoFormer (49.56 km) is substantially lower than that of Informer (76.68 km) and Autoformer (70.86 km), indicating better spatial generalization. The improvements persist on the 2024 test year, confirming that TyphoFormer generalizes well to recent, unseen typhoon events. These results highlight the effectiveness of our method in leveraging language-augmented prior knowledge for robust typhoon trajectory forecasting.
3.2. Case study
To assess the practical forecasting capability of TyphoFormer, we present a detailed case study on Hurricane MILTON. As shown in Figure 2, it developed in the Gulf of Mexico and made landfall in Florida in 2024. Notably, TyphoFormer achieves superior spatial fidelity across the entire storm path, especially in three critical segments where many baselines exhibit significant deviations: (1) Early recurvature in the Gulf of Mexico: In the initial west-to-east drift across the central Gulf, TyphoFormer correctly anticipates the cyclone’s gentle northward bend around longitude 90°W, closely aligning with the observed dynamics. In contrast, Informer and CLIPER incorrectly maintain a more zonal trajectory, overshooting westward before eventually turning north. This suggests that TyphoFormer benefits from prompt-informed inductive priors that enhance recognition of subtle early steering shifts. (2) Florida landfall and curvature east of 85°W: The most critical test of trajectory fidelity occurs near the landfall point. TyphoFormer precisely captures both the timing and curvature of the landfall arc over Florida’s western coast. In contrast, LSTM and GRU underestimate the curvature, producing flatter tracks that result in eastward positional bias post-landfall. These models tend to smooth transitions excessively, failing to model the nonlinear influence of land-sea friction and subtropical ridge interactions. (3) Post-landfall drift into the Atlantic: After crossing the Florida peninsula, TyphoFormer continues to trace the observed northeastward drift offshore with minimal divergence. Informer and Autoformer begin to diverge more drastically in this segment, due to error accumulation and limited ability to represent mesoscale dynamics once the storm enters the open Atlantic. TyphoFormer’s attention-based structure with fused language conditioning contributes to its resilience in such dynamically uncertain regimes.

### v3

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: The 33rd ACM International Conference on Advances in Geographic Information Systems; November 3–6, 2025; Minneapolis, MN, USAbooktitle: The 33rd ACM International Conference on Advances in Geographic Information Systems (SIGSPATIAL ’25), November 3–6, 2025, Minneapolis, MN, USAdoi: 10.1145/3748636.3763223isbn: 979-8-4007-2086-4/2025/11ccs: Information systems Data mining
@@ 1. Introduction
Typhoons and hurricanes pose significant threats to coastal regions, causing devastating damage to infrastructure, ecosystems, and human lives (ramachandran2025addressing; wachnicka2020major). The southeastern United States, especially Florida, is facing increasingly frequent and intense storms (oliver2019climate; malmstadt2009florida). For instance, Hurricane IAN in 2022 resulted in over 150 fatalities and an estimated \$113 billion in economic losses, making it one of the costliest hurricanes in U.S. history (wang2024impacts; bushong2023critical). With the rise of sea surface temperatures and the evolution of climate patterns, accurate typhoon trajectory forecasting has become more critical than ever for emergency planning, resource allocation, and early evacuation (wang2024changes). Reliable predictions of storm movement not only support disaster mitigation efforts but also enable authorities to issue timely warnings, minimizing potential casualties and disruption (yu2022impact). In light of the substantial socio-economic impact of such extreme weather events, there is a growing demand for advanced, data-driven forecasting methods that can complement or enhance traditional meteorological models.
@@
Existing approaches for typhoon track forecasting generally fall into two main categories: (I) time series modeling methods and (II) physics-based numerical models. To address the need for data-driven forecasting, a growing body of work has explored modeling typhoon tracks as temporal sequences using deep learning techniques (wang2024multimodal). These methods treat the storm’s historical positions and attributes as a time series and aim to predict its future trajectory based on learned temporal patterns using deep learning-based methods. Early approaches employed recurrent neural network (RNN)-based architectures such as LSTM, GRU, and their variants to capture sequential dependencies in storm movement (song2022novel; wei2020nearshore). More recent advances have leveraged Transformer-based architectures, which offer improved long-range modeling capabilities and greater scalability for sequence forecasting (zhou2021informer; wu2021autoformer). Methods such as InFormer (zhou2021informer) and AutoFormer (wu2021autoformer) have demonstrated promising results in related time series domains, particularly for long-horizon prediction tasks. While these models excel in learning from past trajectory data, they lack an awareness of external contextual information, such as meteorological semantics or high-level event dynamics, that may further enhance predictive accuracy.
@@
As a basic tool in traditional meteorological analysis, physics-based numerical models are also widely used in operational typhoon forecasting (iwasaki1987performance; knaff2007statistical). Numerical models simulate atmospheric dynamics by solving complex partial differential equations derived from physical laws (bauer2015quiet), such as the Navier–Stokes Equations (rozanova2010typhoon) and Boundary Layer Structure Modeling (kepert2010slab). Prominent forecasting systems, including the the Global Forecast System (GFS) (halperin2020comparison) and Climatology and Persistence Model (knaff2007statistical), use large-scale initial conditions, such as pressure fields, wind vectors, and sea surface temperatures, to estimate future storm trajectories. While these models offer physical interpretability and have been refined through decades of meteorological research, they require gigantic computational resources (e.g., supercomputing center) and are sensitive to initial condition uncertainties. Furthermore, their performance may degrade in data-sparse regions or during rapidly evolving storm phases, where high-resolution input data are limited or delayed. As a result, complementary deep learning and data-driven techniques have been increasingly explored to address these limitations inherent in traditional approaches.
@@ 2.2. Language Contextual Prompt Generation
To enhance the model’s contextual understanding of meteorological dynamics, we generate natural language descriptions (referred to as semantic prompts) for each time step in the trajectory data. Unlike conventional approaches that rely solely on structured numerical inputs, our method introduces a textual layer of information that captures the same data in a human-readable, semantically coherent format. These prompts are automatically generated using a GPT-4o through API-based batch inference. Given a single line record from the trajectory dataset, such as:
AL142024_MILTON,20241010,0030,L,HU,27.4N,82.6W,100,
958,180,170,110,220,60,60,70,90,30,30,30,30,20
@@
These natural language prompts encode meteorological knowledge in a compact form that would otherwise require explicit domain understanding. By providing a semantic summary of each state, they serve as high-level guidance to assist the forecasting model in interpreting patterns and trends within the trajectory. All prompts are generated offline using the GPT-4o API and cached for training and evaluation. This step introduces minimal computational overhead and significantly enriches the input representation.
@@ 2.3. Text Embedding and Feature Fusion
[…]
@@ 3. Experiments
Dataset. Experiments are carried on the well-known HURDAT2 dataset, which is the official North Atlantic hurricane database maintained by the U.S. National Hurricane Center. It provides detailed 6-hour records of tropical cyclones, including 22 critical meteorological features. Each typhoon event is recorded as a chronological sequence. We use the records from year 2004 to 2021 for training, while the data samples from 2022 to 2024 are used for testing.
Evaluation metrics. We use the commonly adopted prediction criteria Mean Absolute Error (MAE) and a task-specific criteria–Spherical Distance Error ($\Delta\text{R}$). $\Delta\text{R}$ measures the geospatial distance between the prediction and the ground-truth trajectories, which is calculated as: $\Delta\text{R}=\Delta\text{Re}\cdot\text{arccos}(\sin\varphi_{p}\sin\varphi_{r}+\cos\varphi_{p}\cos\varphi_{r}\cos(\lambda_{p}-\lambda_{r}))$, where $\text{Re}=6371km$ is the Earth’s radius, $(\varphi_{p},\lambda_{p})$ and $(\varphi_{r},\lambda_{r})$ are the predicted / ground-truth latitude and longitude, respectively.
Baseline methods. We compare our TyphoFormer [^2] https://github.com/LabRAI/TyphoFormer against two categories of baselines. The first category is the classical numerical equation-based models, we choose the representative (I) CLIPER (Climatology and Persistence Model) (knaff2007statistical), which extrapolates storm trajectories using historical patterns and persistence assumptions. The second category is time-series forecasting models, including recurrent models like (III) LSTM and (IV) GRU, as well as Transformer-based or MLP-based advanced architectures like (V) Informer (zhou2021informer), (VI) Autoformer (wu2021autoformer), and (VII) TSMixer (ekambaram2023tsmixer). These baselines provide a diverse comparison across classical statistical models and advanced deep learning frameworks.
@@ 3.1. Main Results
The best values are in bold font, the second-best values are indicated with underlining.
[…]
6h 12h 18h 24h 6h 12h 18h 24h
CLIPER 0.235 0.275 0.310 0.368 34.265 42.205 51.632 58.268
GRU 0.367 0.405 0.493 0.640 50.480 69.397 90.875 103.894
LSTM 0.392 0.431 0.583 0.828 46.096 71.365 95.412 112.663
Informer 0.289 0.318 0.392 0.483 37.592 46.435 56.433 76.684
Autoformer 0.263 0.286 0.357 0.433 39.836 47.183 63.775 70.862
TSMixer 0.214 0.268 0.297 0.353 35.720 45.265 50.330 62.910
TyphoFormer 0.188 0.242 0.261 0.312 31.539 38.084 42.435 49.562
[…]
6h 12h 18h 24h 6h 12h 18h 24h
CLIPER 0.221 0.269 0.294 0.361 33.915 44.542 48.580 57.642
GRU 0.359 0.394 0.491 0.632 49.674 70.357 92.542 105.114
LSTM 0.388 0.422 0.587 0.835 44.203 69.051 94.555 109.108
Informer 0.290 0.316 0.395 0.487 37.843 48.791 57.445 78.497
Autoformer 0.257 0.284 0.358 0.435 39.128 46.085 65.039 72.827
TSMixer 0.228 0.265 0.305 0.347 35.148 43.935 52.501 63.118
TyphoFormer 0.185 0.241 0.263 0.317 31.274 37.934 42.973 50.881
@@
[…]
3.2. Case study on Hurricane MILTON
Hurricane MILTON is developed in the Gulf of Mexico and made landfall in Florida in 2024 (see Figure 2). Notably, TyphoFormer achieves superior spatial fidelity across the entire storm path, especially in three critical segments where many baselines exhibit significant deviations: (1) Early recurvature in the Gulf of Mexico: In the initial west-to-east drift across the central Gulf, TyphoFormer correctly anticipates the cyclone’s gentle northward bend around longitude 90°W, closely aligning with the observed dynamics. In contrast, Informer and CLIPER incorrectly maintain a more zonal trajectory, overshooting westward before eventually turning north. This suggests that TyphoFormer benefits from prompt-informed inductive priors that enhance recognition of subtle early steering shifts. (2) Florida landfall and curvature east of 85°W: The most critical test of trajectory fidelity occurs near the landfall point. TyphoFormer precisely captures both the timing and curvature of the landfall arc over Florida’s western coast. In contrast, LSTM and GRU underestimate the curvature, producing flatter tracks that result in eastward positional bias post-landfall. These models tend to smooth transitions excessively, failing to model the nonlinear influence of land-sea friction and subtropical ridge interactions. (3) Post-landfall drift into the Atlantic: After crossing the Florida peninsula, TyphoFormer continues to trace the observed northeastward drift offshore with minimal divergence. Informer and Autoformer begin to diverge more drastically in this segment, due to error accumulation and limited ability to represent mesoscale dynamics once the storm enters the open Atlantic. TyphoFormer’s attention-based structure with fused language conditioning contributes to its resilience in such dynamically uncertain regimes.
@@ 4. Conclusions
Acknowledgements.
This work is supported in part by the start up grant and the FYAP grant program provided by Florida State University.